OpenStudy (anonymous):

Factor the equation 4x^2-20x=0

OpenStudy (anonymous):

please show work

OpenStudy (unklerhaukus):

\[4x^2-20x=0\]\[(4)(x)(x)-(4)(5)(x)=0\]\[=4x(x-5)=0\] \[x=0,5\]

OpenStudy (anonymous):

yup answer is x=0 x=5 4x(x-5)=0 x=0,x=5

OpenStudy (anonymous):

@rickybobby13 gets??

OpenStudy (anonymous):

I still I'm having a hard time understanding this.

OpenStudy (unklerhaukus):

which step do you not fully understand

OpenStudy (anonymous):

how to go from \[(4)(x)(x)-(4)(5)(x)=0\] to \[4x(x-5)=0\]

OpenStudy (anonymous):

he isolate the 4x because it is common

OpenStudy (anonymous):